OpenStudy (anonymous):

Are a rational expression and its simplified forms equivalent? Provide mathematical examples to support your opinions. You may use tables or graphs to organize and present your thoughts.

OpenStudy (anonymous):

@ImPooka I need help on what you mean by rational expression. Are we talking 16/12 is equivalent to 4/3? Or, are we wondering if y =[(x-1)(x+2)] /(x+2) is equivalent to y = (x - 1) ?

OpenStudy (anonymous):

16/12 equivalent to 4/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, thanks chief

OpenStudy (anonymous):

np